Project

General

Profile

2015-11-17 memo-physics-report

2015/03

  • Third draft run plan document is at http://micewww.pp.rl.ac.uk/projects/analysis/documents
  • Lots of uncertainty around hardware, so we shall see what happens (ops meeting later)
  • FC alignment:-
    • Rogers has looked at settings, ~ 50 A solenoid mode looks okay (or 70 A flip mode); still not finalised
    • Sophie is back from vacation and will start looking through "reconstruction" i.e. can we get good enough resolutions with TOF cut?
  • Q789 switched off for TOF0 alignment
    • nb this is a short run but should be done for TOF0 alignment
  • SSU switched on with ECE:-
    • Uncertainty around the hardware
    • VB looking at trim coil currents
    • Nb: needs some modelling of iron here - feeding in Jo Langlands (Sheffield grad student) to this effort + field mapping
  • Straight tracks
    • Staring at MuScat paper, looks like they were systematics dominated (errors ~ 50k muons)
    • Working on the statistics analysis (can we get 50 k muons within momentum cut and radial cut at the absorber? what should the cuts be?)

3/4 December Review

  • Step IV
    • asked Ao to push through a new round of optics studies - he was distracted by PASI last week
    • will then need to process at least one setting with full reconstruction
  • Demonstration of ionisation cooling
    • Baseline and baseline w/o M1 is at a reasonable point, using RMS emittance analysis
    • CR is looking at generic phase space volume analyses (not RMS emittance); to fight the "do nothing" crowd
    • Jaroslaw is looking at using a third FC as a matching element; he didn't find a match at 4 T only at 2 T

January optics Review

  • v1 agenda handed to KL, v2 attached (some familiar names)
  • await revised ToR - but we need to start the writing up process now to have a hope to converge by January; agenda becomes the latex skeleton
  • I am scared by the timescale

Step IV

  • Precision measurement of emittance
    • VB working through the "TOF tracks" analysis (maybe she is in the meeting?), turning it into a reconstruction routine
    • Awaits tracker reconstruction/geometry fixes
  • Tracker alignment
    • Discussion with Melissa/Jason/Durga/Tony Millington
    • John Palin et al aligned trackers to monuments on the end face of SSU/SSD
    • Tony Millington et al surveyed monuments on the outer rim of SSU/SSD
    • end-face monuments are now obscured
    • Jason Tarrant is in contact with John Palin to recover any (raw) survey data he may have, hand off to Tony
    • Tony Millington will make a remedial survey of SSU on Wednesday this week when the diffuser goes in
    • remedial survey on SSD may not be possible due to interference from cables and pipes

Cooling demo

  • No update on the paper, chasing on Jaroslaw "cooling demo with focus coil" optics
  • No update on the tolerances study, will chase Alan

optics-review-agenda-v2.xls (9 KB) Rogers, Chris, 17 November 2015 12:52